Hagit's suggestion is really good. If there's a hot spot on the filament, the electrons could be shot off to the side, so to speak, directly into the "box" and never make it to the trap.
I'm not exactly sure what happens if the trap and box short. I'm guessing you won't pick it up when you check, but the trap can twist, and if there's only a tiny gap between them--you might get shorting at high voltage and with some gas pressure in the source.
The same can be said for the hood on the filament, and Chris is probably on to something in this case. Also, if the magnet is not completely secure, it might sag and the holder get a little too close to the filament housing for comfort. Over tightening might lead to the same problem--not sagging, but magnet holder getting too close to the filament housing. Make sure the spacers are not flaring.
I've also had issues with the filament holder not sitting flat against the box. If you're using washers there, make sure they're only on the alumna, and not catching on the lip of cutout in the box. That can prevent the filament holder from sitting flat against the box.
On 12/20/2014 3:14 PM, Herbert Tobias wrote:
Still having high box current problem on MAT 253.
As per Gerard's suggestion, I checked the trap lead connection. When removed source, it was good and connected.As per Hagit, I already tried reattaching filament.Today I replaced the filament with a new one (actually SIS repaired) and now Box = 1.24, Trap = 0.26 mA. Just arced and now Box =1.38, Trap = 0.12 mA.
Yesterday with previous filament (a new Thermo one), when I heat up source to ~60C, the Box = 0.65, Trap = 0.85 with OK signal (40% less than ideal).But then went back to Box = 1.5 mA, Trap = 0 mA after source got hotter (80C). And stayed as such when cooled. Was arcing often.
Any other suggestions?
On Fri, Dec 19, 2014 at 11:34 AM, hagit affek <[log in to unmask]> wrote:
This happened to us once before. the trapped hovered close to zero instead of our usual ~0.5 and box was close to 1.5. signal intensity was normal but the delta values were very noisy and probably the beam was not stable. the solution was to open the source again, detach the filament and then reattach it.
Hagit P. Affek
Dept. of Geology & Geophysics
210 Whitney Ave.
PO Box 208109
New Haven, CT 06520-8109
On 12/19/14, 11:11 AM, Herbert Tobias wrote:
This week I cleaned the source of our MAT 253 and installed a new Thermo filament. Apart from resolving a bunch of annoying issues over the days, and re-cleaning yesterday, the Trap current on the system is currently reading Zero (0), while the Box current reads the emission I set in the focus settings (i.e. 1.5 mA or 1.0 mA etc). All through the week and just this morning the Box and Trap current was reading approximately equal (i.e. 0.75 mA and 0.75 mA). Right now for the past couple hours it is lopsided as noted above (Box= 1.5 mA and Trap = 0 mA). But the instrument does provide strong signal for CO2 (m/z 44 = ~4500 mV at 1.5 bar on Ref 1 of a Conflo III). I measure no shorts to the Trap. I decided to start baking the system for awhile and see what happens. Any thoughts?
(Title of e-mail sounds like a psychological thriller, eh?)
thanks and regards
Dr. Herbert Tobias
Sr. Research Associate
Division of Nutritional Sciences
244 Garden Ave
Ithaca NY 14853
Dr. Herbert TobiasSr. Research AssociateDivision of Nutritional Sciences244 Garden AveCornell UniversityIthaca NY 14853